# PKF: Probabilistic Data Association Kalman Filter for Multi-Object Tracking

### v1

@@ 4.3.2 Weight computation
where $\mathcal{D}_{t}^{\Phi_{t}}$ is the set of association events that have $\Phi_{t}$ false measurements. The probability of measurement $\mathbf{z}_{t,k}$ being assigned to object $\mathbf{x}_{t,j}$ is computed in [2, Eq. 51] as
@@ 4.3.3 Tracking comparison
Figure 2: Object tracking simulation showing ground-truth object trajectories (dashed lines), noisy object position measurements (green dots), and clutter measurements (yellow crosses).
@@
Using the same association weight computation, we compare the performance of JPDAF [2] and our PKF with 2D simulated data. In the simulation, each object is a 2D point and moves along an 8-shape trajectory. The detection probability of each object is $p^{t}_{D}=0.9$, the measurement noise is a zero-mean Gaussian with covariance $\operatorname*{diag}(0.75,0.75)$, and clutter is sampled uniformly in the range of $[-10,10]^{2}$ around each ground-truth point. A visualization of 3 simulated objects can be found in Figure 2.
The errors and speed of tracking 3 and 5 objects in simulation are shown in Table 1. PKF achieves lower tracking errors than JPDAF [2] implemented by [19]. Our intuition as to why PKF achieves lower errors is provided in the first paragraph of Sec. 4.3.1. The experiments also show that PKF can track at a comparable speed with JPDAF.
@@
We also compare the two algorithms with a larger number of objects, i.e., 10 objects, under different levels of measurement noises. In this setting, we increase the detection probability to $0.95$ to avoid frequent tracking failures. The clutter generation remains the same and we test with measurement noises from $0.2$ to $0.75$. The results can be found in Figure 3. We can see that our filter obtains lower tracking errors in most cases. The number of failed tracks is almost the same as JPDAF [2]. The average fps of JPDAF with 10 objects is $1284$ and that of ours is $1080$.
Table 1: Comparison of tracking error ($l2$ norm in meters) and speed (frames per second (fps)) among binary association, JPDAF [2], and PKF in simulation.
$N_{obj}$ Method Obj 1 Obj 2 Obj 3 Obj 4 Obj 5 Avg fps
3 Binary 2.16 1.36 1.80 - - 1.78 5062
JPDAF [2] 0.70 0.60 0.65 - - 0.65 3631
PKF 0.70 0.57 0.58 - - 0.62 3235
5 Binary 8.37 19.36 17.34 11.37 14.53 14.20 3061
JPDAF [2] 0.68 0.73 0.62 0.63 0.66 0.66 2567
PKF 0.60 0.68 0.58 0.61 0.61 0.62 2172
@@
Figure 3: Ablation study with 10 objects in the simulation under different noise scales. Left: tracking errors in meters. Right: number of failed tracks, i.e. those with tracking errors larger than $5$.
Figure 4: Illustration of our hybrid data association. The detections and tracked objects are shown in green and orange, respectively. The first row shows the result of binary association via the Jonker-Volgenant algorithm [21]. The second row is our proposed hybrid data association. In the first image, objects 1 and 3 are moving toward each other, as indicated by the red arrows. In the second image, the detections overlap highly and our ambiguity check recognizes these as ambiguous (inside the dashed red boxes). The binary association causes an ID switch but our method can track the objects correctly, as shown in the third image (inside the dashed red boxes).
5 Application to multi-object tracking
In this section, we apply our PKF to the MOT task.
5.1 Algorithm design
To make the filter work efficiently in practice, we introduce a hybrid data association scheme that applies probabilistic data association only to a subset of ambiguous measurements. Also, we force the state covariance matrix to be block diagonal so that each track is estimated by a separate filter, although the states are still considered jointly when evaluating the data association probabilities. We follow SORT [6] and model each object’s state as $\mathbf{x}=[u,v,s,r,\dot{u},\dot{v},\dot{s}]^{\top}$, where $u$ and $v$ represent the horizontal and vertical pixel location of the center of the object bounding box, while $s$ and $r$ represent the scale (area) and the aspect ratio of the object’s bounding box respectively. The detections are converted into $\mathbf{z}=[u,v,s,r]^{\top}$.
Hybrid data association Constructing a probability matrix using all measurements and states and computing the association weights according to Proposition 1 can be slow and potentially harm the tracking accuracy due to low-probability measurements. A binary matching algorithm, such as the Hungarian [23] or the Jonker-Volgenant [21], is efficient but may make incorrect hard decisions when there is ambiguity. We propose a hybrid association scheme to take advantage of both matching types. Given measurements $\mathbf{z}_{t,k}$, $k=1,\ldots,M_{t}$ and states $\mathbf{x}_{t,j}$, $j=1,\ldots,N$, we first compute a score matrix $S_{t}\in\mathbb{R}^{M_{t}\times N}$ with entries $[S_{t}]_{kj}$ equal to the intersection over union (IoU) for each measurement-object bounding-box pair. A binary matching algorithm is applied to $S_{t}$ to obtain a binary matching weight matrix. Then, we perform an ambiguity check. For each row $k$ of $S_{t}$ (corresponding to measurement $\mathbf{z}_{t,k}$), we sort the scores from high to low. Supposing the rank is $\mathbf{x}_{t,j_{1}}\ \mathbf{x}_{t,j_{2}}\cdots\mathbf{x}_{t,j_{N}}$, if object $\mathbf{x}_{t,j_{2}}$ has a score over a threshold $\tau_{ambig}$, e.g., $90\%$ the score of object $\mathbf{x}_{t,j_{1}}$, then $\mathbf{x}_{t,j_{1}}$ and $\mathbf{x}_{t,j_{2}}$ are ambiguous objects, and the measurement $\mathbf{z}_{t,k}$ is an ambiguous measurement. Then, we compare $\mathbf{x}_{t,j_{2}}$ and $\mathbf{x}_{t,j_{3}}$ and so on. We repeat this for all objects and measurements to obtain an ambiguous set. Measurements and objects with their matches marked as ambiguous will also be added to the ambiguous set. Proposition 1 is only applied to the ambiguous set of measurements and objects. Finally, we replace the elements in the initial binary matching weight matrix with the corresponding probabilistic weights. An illustration of how our proposed probabilistic data association Kalman filter benefits the ambiguous case is shown in Figure 4 where binary association causes a wrong ID switch while our proposed association can obtain a correct ID.
Each object as a single filter Instead of storing all objects in the filter state, we create a filter for each new object (which is equivalent to enforcing a block-diagonal covariance matrix). This avoids a large state covariance and observation Jacobian $\bar{H}_{t+1}$ in (17), which speeds up the tracking especially when the number of objects is large. After we get the association weight matrix $W$, each column vector $\mathbf{w}_{j}$ corresponds to a tracked object with measurement association weights. All measurements with a weight larger than a threshold $\tau_{weight}$ are used to update the tracker by Proposition 2. The thresholding helps exclude potential outliers and achieve stable computation.
5.2 Evaluation
Datasets
We evaluate PKF on multiple MOT datasets, including MOT17 [27], MOT20 [15], and DanceTrack [40]. MOT17 and MOT20 focus on crowded pedestrian tracking in public places. DanceTrack is a dancing-scene dataset, where the objects have similar appearances, move fast, highly non-linearly, and cross over each other frequently, thus imposing a higher requirement on data association.
Metrics We use higher order tracking accuracy (HOTA) [24] as our main metric because it balances between detection accuracy and association accuracy. We also report the AssA [24] and IDF1 [38] metrics which emphasize data association accuracy and the MOTA [5] metric which emphasizes the detection accuracy.
@@ 5.2 Evaluation
Table 2: Results on MOT17 [27] testset with private detections. We share detections with ByteTrack [47] and OC-SORT [13].
@@
5.3 Benchmark results
MOT17 and MOT20
The quantitative results on MOT17 [27] and MOT20 [15] are shown in Table 2 and 3. To get a fair comparison, we use the same detections as ByteTrack [47] and OC-SORT [13] and inherit the linear interpolation. While ByteTrack [47] does multiple rounds of associations and OC-SORT makes use of velocities, does object recovery, and re-updates, PKF only does one round of association using only bounding boxes and performs one single update step. We also do not use the adaptive detection thresholds as in ByteTrack [47] for generalization. We can see that our method is able to achieve comparable results with those two methods [47, 13] and is better in data association as indicated by the higher IDF1 and AssA and achieves higher HOTA results. Note that our method is compatible with techniques used by other methods, like multiple rounds of associations [47], taking into account velocities [13], and doing object recovery [13] but we did not incorporate these to keep the results clear.
Table 3: Results on MOT20 [15] testset with private detections. We share detections with ByteTrack [47] and OC-SORT [13].
@@ 5.3 Benchmark results
Table 4: Results on DanceTrack [40] testset. Methods in blue share detections.
@@
PKF 51.7 79.8 33.5 89.1 50.1
PKF + OCM&OCR 55.4 79.5 38.7 89.2 54.9
DanceTrack The results on DanceTrack [40] are shown in Table 4. We can see that PKF achieves better results than SORT [6] and ByteTrack [47]. Since the movement is significant and the number of objects is usually fixed in a dancing scene, we find that OCM (velocity) and OCR (object recovery) proposed in OC-SORT [13] are useful. By incorporating these two components, we can further outperform OC-SORT without ORU (re-update) [13]. In this dataset, there are a lot of occlusions and the dancers can move dramatically. The increase of AssA and IDF1 shows again the advantage of PKF in terms of association. We can see that PKF obtains lower DetA than [13]. This is because the probabilistic data association can incorporate incorrect detections in the update step, making the position estimates slightly worse. However, we can still obtain higher HOTA since $\text{HOTA}_{\alpha}$ at a certain threshold $\alpha$ can be decomposed into two parts, a detection accuracy score (DetA) and an association accuracy score (AssA):
$\text{HOTA}_{\alpha}=\sqrt{\text{DetA}_{\alpha}\cdot\text{AssA}_{\alpha}}.$ (20)
The increased HOTA shows the increased association quality can compensate for relatively worse position updates.
Table 5: Ablation study of ambiguity check (A.C.) and weight thresholding (W.T.) on the validation sets of MOT17 [27] and DanceTrack [40] in terms of HOTA [24], IDF1 [38], and average tracking time per frame ($t$).
@@
PKF w/o W.T. 66.6 78.1 6.3 52.0 50.4 3.8
PKF complete 66.8 78.5 6.2 53.3 53.2 3.7
Ablation study We study the effect of the ambiguity check and weight thresholding on the validation sets of MOT17 and DanceTrack. The results are shown in Table 5. We observed that associating all measurements and tracks directly not only decreases the association quality, implied by the low HOTA and IDF1, but also makes the tracking slow. Adding weight thresholding can further improve the results as low-probability measurements are usually wrong. Weight thresholding also reduces the time a bit as we use fewer measurements in the update step. The complete PKF algorithm obtains better results than binary association with almost the same tracking speed.
6 Conclusion
We derived a new formulation of the Kalman filter with probabilistic data association by formulating a variational inference problem and introducing data association as a latent variable in the EM algorithm. We showed that, in the E-step, the association probabilities can be computed as matrix permanents, while the M-step led to the usual Kalman filter prediction and update steps but with an extended measurement vector and noise covariance matrix capturing possible data associations. Our experiments demonstrated that our filter can outperform the JPDAF and can achieve top-10 results on MOT benchmarks even without using deep features. Our algorithm is not restricted to the MOT application and can serve as a general method for estimation problems with measurement ambiguity.
@@ 6 Conclusion
7 Proof of Proposition 1
Define sets $\mathcal{K}^{-k}_{t}:=\{1,...,M_{t}\}\backslash\{k\}$, $\mathcal{J}^{-j}:=\{1,...,N\}\backslash\{j\}$, and a function $\bar{\delta}_{t}:\mathcal{K}^{-k}_{t}\rightarrow\mathcal{J}^{-j}$. Then, (6) can be rewritten as:
@@ 8 Proof of Proposition 2
$=\begin{bmatrix}\Sigma_{t}^{-1}\boldsymbol{\mu}_{t}-F^{\top}W^{-1}G\mathbf{u}_{t}\\ W^{-1}G\mathbf{u}_{t}+\bar{H}_{t+1}^{\top}\bar{V}_{t+1}^{-1}\bar{\mathbf{z}}_{t+1}\end{bmatrix}$
@@
$=\begin{bmatrix}\Sigma^{-1}_{t}\boldsymbol{\mu}_{t}-F^{\top}W^{-1}G\mathbf{u}_{t}\\ \uwave{(\Sigma^{+}_{t+1})^{-1}(F\boldsymbol{\mu}_{t}+G\mathbf{u}_{t})+\bar{H}^{\top}_{t+1}\bar{V}_{t+1}^{-1}\bar{\mathbf{z}}_{t+1}}\end{bmatrix}$

### v2

@@ IV-C2 Comparison with JPDAF
where $\mathcal{D}_{t}^{\Phi_{t}}$ is the set of association events that have $\Phi_{t}$ false measurements. The probability of measurement $\mathbf{z}_{t,k}$ being assigned to object $\mathbf{x}_{t,j}$ is computed in [10, Eq. 51] as
@@ IV-C3 Tracking error
Figure 2: Object tracking simulation showing ground-truth object trajectories (dashed lines), object measurements (green dots), and clutter (false) measurements (yellow crosses).
@@
We compare our PKF to PMHT [9] and JPDAF [10] with 2D simulated data, where each object is a point and moves along an 8-shape trajectory. The detection probability of each object is $p^{t}_{D}=0.9$, the measurement noise is a zero-mean Gaussian with covariance $\operatorname*{diag}(0.75,0.75)$, and clutter is sampled uniformly in the range of $10,10]^{2}$ around each ground-truth point. We use $\lambda=0.125$ for the spatial density of the Poisson pmf. A visualization of 3 simulated objects can be found in Figure 2.
@@
The errors of tracking 3 and 5 objects in simulation are shown in Table I. PMHT [9] works slightly better than binary ssociation but much worse than PKF and JPDAF [10] due to the non-exclusive data association. PKF achieves lower tracking errors than JPDAF implemented by [42]. Our intuition as to why PKF achieves lower errors is provided in the first paragraph of Sec. IV-C2. We also compare PKF to JPDAF under different levels of measurement noise with 10 objects. In this setting, we increase the detection probability to $0.95$ to avoid frequent tracking failures. The clutter generation remains the same, and we test with measurement noises from $0.2$ to $0.75$. The results can be found in Figure 3. We can see that our filter obtains lower tracking errors in most cases. The number of failed tracks is almost the same as JPDAF [10].
IV-C4 Update speed
We compare the update speeds of these methods with different numbers of objects. The results can be found in Table II. The experiments show that PKF can track at a comparable speed with JPDAF.
TABLE I: Comparison of tracking errors ($l2$ norm in meters).
$N_{obj}$ Method Obj 1 Obj 2 Obj 3 Obj 4 Obj 5 Avg
3 Binary 2.16 1.36 1.80 - - 1.78
PMHT [9] 1.50 1.29 1.55 - - 1.45
JPDAF [10] 0.70 0.60 0.65 - - 0.65
PKF 0.70 0.57 0.58 - - 0.62
5 Binary 8.37 19.36 17.34 11.37 14.53 14.20
PMHT [9] 12.41 12.34 12.43 12.10 11.21 12.10
JPDAF [10] 0.68 0.73 0.62 0.63 0.66 0.66
PKF 0.60 0.68 0.58 0.61 0.61 0.62
TABLE II: Update time (ms) per frame given association weights.
Number of objects 3 5 10 20 Avg
Vanilla Kalman filter 0.08 0.12 0.20 0.36 0.19
PMHT [9] 0.11 0.16 0.28 0.55 0.28
JPDAF [10] 0.28 0.39 0.78 1.47 0.73
PKF 0.31 0.46 0.93 1.76 0.87
@@ IV-C4 Update speed
Figure 3: Ablation study with 10 objects in the simulation under different noise scales. Left: tracking errors in meters. Right: number of failed tracks, i.e., those with tracking errors larger than 55.
@@
TABLE III: Results on MOT17 [14] testset with private detections. Methods in blue share detections.
@@
TABLE IV: Results on MOT20 [15] testset with private detections. Methods in blue share detections.
@@
V Application to multi-object tracking
In this section, we apply our PKF to the MOT task.
V-A Algorithm design
[…]
Ambiguity check Gating is very important in Kalman filtering [10]. We propose a simple yet efficient way to find ambiguous measurements in the meantime of removing potential outliers (gating). Given measurements $\mathbf{z}_{t,k}$, $k=1,\ldots,M_{t}$ and states $\mathbf{x}_{t,j}$, $j=1,\ldots,N$, we first compute a score matrix $S_{t}\in\mathbb{R}^{M_{t}\times N}$ with entries $S_{t}]_{kj}$ equal to the intersection over union (IoU) for each measurement-object bounding-box pair. Then, we perform an ambiguity check. For each row $k$ of $S_{t}$ (corresponding to measurement $\mathbf{z}_{t,k}$), we sort the scores from high to low. Supposing the rank is $\mathbf{x}_{t,j_{1}}\ \mathbf{x}_{t,j_{2}}\cdots\mathbf{x}_{t,j_{N}}$, if object $\mathbf{x}_{t,j_{2}}$ has a score over a threshold $\tau_{ambig}$, e.g., $90\%$ the score of object $\mathbf{x}_{t,j_{1}}$, then $\mathbf{x}_{t,j_{1}}$ and $\mathbf{x}_{t,j_{2}}$ are ambiguous objects, and the measurement $\mathbf{z}_{t,k}$ is an ambiguous measurement. Then, we compare $\mathbf{x}_{t,j_{2}}$ and $\mathbf{x}_{t,j_{3}}$ and so on. We repeat this for all objects and measurements to obtain an ambiguous set. Measurements and objects with their matches marked as ambiguous will also be added to the ambiguous set. Proposition 1 is only applied to the ambiguous set of measurements and objects. The rest of the measurements and objects are associated with binary association [31, 48]. An illustration of how our proposed probabilistic data association Kalman filter benefits the ambiguous case is shown in Figure 4, where binary association causes a wrong ID switch while our proposed association can obtain a correct ID.
We create a filter for each new object (equivalent to enforcing a block-diagonal covariance matrix). This avoids a large state covariance and observation Jacobian $\bar{H}_{t+1}$ in (17), which speeds up the tracking especially when the number of objects is large. We still compute the association weight matrix $W$ jointly, each column vector $\mathbf{w}_{j}$ corresponds to a tracked object with measurement association weights. Measurements with a weight larger than a threshold $\tau_{weight}$ are used to update the tracker by Proposition 2, which helps exclude potential outliers and achieve stable computation.
Figure 4: Illustration of our data association. The detections and tracks are shown in green and orange. The first row shows the result of binary association via [48]. The second row is our proposed data association. In the first image, the detections highly overlap, and our ambiguity check recognizes these as ambiguous (dashed red boxes). The binary association causes an ID switch, but our method can track the objects correctly, as shown in the third image.
V-B Evaluation
Datasets We evaluate PKF on multiple datasets, including MOT17 [14], MOT20 [15], and DanceTrack [11]. MOT17 and MOT20 focus on crowded pedestrian tracking in public places. DanceTrack is a dancing-scene dataset, where the objects have similar appearances, move fast, highly non-linearly, and cross over each other frequently, thus imposing a higher requirement on data association.
Metrics We use higher order tracking accuracy (HOTA) [13] as our main metric because it balances between detection accuracy and association accuracy. We also report the AssA [13] and IDF1 [12] metrics, which emphasize data association accuracy and the MOTA [49] metric, which emphasizes the detection accuracy.
V-C Benchmark results
In this section, we first show that PKF can get comparable results to other methods [2, 3] by only associating bounding boxes, i.e., without using any other techniques or features. We then show that PKF is compatible with more advanced object features can serve as a backbone, same as SORT [1].
MOT17 and MOT20 The quantitative results on MOT17 [14] and MOT20 [15] are shown in Table III and IV. To get a fair comparison, we use the same detections as ByteTrack [2] and OC-SORT [3] and inherit the linear interpolation. While the direct comparison to our PKF is the vanilla Kalman filter, we can achieve comparable results to ByteTrack [2] and OC-SORT [3], which do multiple rounds of associations [2] , make use of velocities [3], and re-updates [3]. We can see that PKF is able to achieve better data association, as indicated by IDF1 and AssA, and achieves slightly higher HOTA results.
TABLE V: DanceTrack [11] results. Blue means shared detections.
@@ V-C Benchmark results
PKF 53.7 79.7 36.2 89.1 53.5
TABLE VI: Application to Hybrid-SORT-ReID [4].
Dataset Tracker HOTA $\uparrow$ AssA $\uparrow$ MOTA $\uparrow$ IDF1 $\uparrow$
MOT20 Hybrid-SORT-ReID [4] 63.9 64.7 76.7 78.4
Hybrid-PKF-ReID 64.3 65.1 76.7 79.0
DanceTrack Hybrid-SORT-ReID [4] 65.5 52.2 91.8 67.2
Hybrid-PKF-ReID 66.4 53.9 91.6 68.7
DanceTrack The results on DanceTrack [11] are shown in Table V. We can see that PKF achieves better results than SORT [1] and ByteTrack [2]. In this dataset, there are a lot of occlusions and the dancers can move dramatically. The increase of AssA and IDF1 over SORT [1] shows again the advantage of PKF in terms of association.
Application to Hybrid-SORT-ReID Besides data association techniques from [2, 3], Hybrid-SORT-ReID [4] further uses detection scores, a height-enhanced IoU, and a ReID module with deep neural features. We show that our PKF is compatible with these features, just as SORT [1] is, by replacing the SORT part (Kalman filter) with PKF. We name the new method Hybrid-PKF-ReID. With the score matrix computed by Hybrid-SORT-ReID using all these features, we compute the measurement probability as $p(\mathbf{z}_{t,k}|\mathbf{x}_{t,j})\propto\exp(-\alpha/s_{k,j})$, where $s_{k,j}$ is the association score of $\mathbf{z}_{t,k}$ and $\mathbf{x}_{t,j}$. We test Hybrid-PKF-ReID on MOT20 and Dancetrack, which have more ambiguous scenarios. Table VI shows that adding probabilistic association improves Hybrid-SORT-ReID’s results. We show some sequences that PKF improves in Fig. 5, which are highly occluded. This suggests that PKF is effective in highly ambiguous cases.
Figure 5: Visualizations of different datasets.
TABLE VII: Ablation study of ambiguity check (A.C.).
@@
PKF 66.8 78.5 6.2 53.5 53.3 3.7
TABLE VIII: Effect of gating and ambiguity check in the JPDAF [10] on the $l2$ tracking error (meters) in simulation.
Method Obj 1 Obj 2 Obj 3 Obj 4 Obj 5 Avg
Binary 8.37 19.36 17.34 11.37 14.53 14.20
JPDAF [10] w/o gating 19.32 18.31 10.60 16.15 11.63 15.21
JPDAF [10] w/ gating 0.68 0.73 0.62 0.63 0.66 0.66
JPDAF [10] w/ A.C. 0.62 1.47 0.62 0.63 0.58 0.79
TABLE IX: Computation time (ms) of matrix permanent and overall association weights with different matrix sizes.
Matrix size $5\times 5$ $10\times 10$ $15\times 15$ $20\times 20$ $30\times 30$
Nijenhuis et al. [7] $0.003$ $0.01$ $2.0$ $26.5$ $31510.9$
Huber et al. [8] $6.5$ $20.1$ $21.7$ $35.4$ $67.8$
Best $0.003$ $0.01$ $2.0$ $26.5$ $67.8$
Weight compute $0.03$ $0.2$ $14.5$ $1073.6$ $1766.6$
V-D Ablation study
We compare the effect of ambiguity checks and gating, and analyze the computation time of the matrix permanent.
Ambiguity check We compare the performance of PKF with and without ambiguity check in Table VII. We observe that associating all measurements and tracks directly degrades the performance due to updating with low-weight measurements, which are mostly outliers. Introducing the ambiguity check in PKF to discard low-probability measurements before computing the matrix permanents, leads to better results than binary association with almost the same tracking speed. Table VIII shows that our ambiguity check has a similar effect to Chi-squared gating in the JPDAF [10] with 5 objects in simulation. We do not do gating in real-world data since $1/IoU$ is ill-conditioned for Chi-squared test while the ambiguity check is more flexible with different features.
Matrix permanent computation time We analyze the computation time of the matrix permanent used to obtained the association weights. We take the faster matrix permanent method between [7] and [8] (switch at size $20\times 20$). Table IX shows that the speed is very fast for matrices smaller than $15\times 15$ and computations on matrices up to $30\times 30$ run at over $0.5$ fps.
Discussion We observe that achieving effective use of probabilistic data association requires a balance between removing outliers (gating) and keeping in ambiguous measurements. If there is no gating, the outliers will bring bias to the measurements. On the other hand, if only the most likely measurements are considered, wrong hard decisions could harm the estimation. The probability and gating model can be improved to unlock further potential of PKF.
VI Conclusion
We derived a new formulation of the Kalman filter with probabilistic data association by formulating a variational inference problem and introducing data association as a latent variable in the EM algorithm. We showed that, in the E-step, the association probabilities can be computed as matrix permanents, while the M-step led to the usual Kalman filter prediction and update steps but with an extended measurement model. Our experiments demonstrated that our filter can outperform the JPDAF and can achieve good results on MOT benchmarks. Our algorithm is not restricted to the MOT application and can serve as a general method for estimation problems with measurement ambiguity.
@@ VI Conclusion
VII Proof of Proposition 1
[…]
@@ VIII Proof of Proposition 2
$=\begin{bmatrix}\Sigma_{t}^{-1}\boldsymbol{\mu}_{t}-F^{\top}W^{-1}G\mathbf{u}_{t}\\ W^{-1}G\mathbf{u}_{t}+\bar{H}_{t+1}^{\top}\bar{V}_{t+1}^{-1}\bar{\mathbf{z}}_{t+1}\end{bmatrix}$ .
@@
$=\begin{bmatrix}\Sigma^{-1}_{t}\boldsymbol{\mu}_{t}-F^{\top}W^{-1}G\mathbf{u}_{t}\\ \uwave{(\Sigma^{+}_{t+1})^{-1}(F\boldsymbol{\mu}_{t}+G\mathbf{u}_{t})+\bar{H}^{\top}_{t+1}\bar{V}_{t+1}^{-1}\bar{\mathbf{z}}_{t+1}}\end{bmatrix}$ .
@@
IX Motion and measurement model matrices
In this section, we provide details for the motion and measurement matrices. For the motion model, since we store the velocities in the state, the motion model is $\mathbf{x}_{t+1}=[u_{t}+\dot{u}_{t},v_{t}+\dot{v}_{t},s_{t}+\dot{s}_{t},r_{t},\dot{u}_{t},\dot{v}_{t},\dot{s}_{t}]$, and there are no control inputs, so we can set $G=0$. We have
$F=I_{7\times 7}+\begin{bmatrix}0_{3\times 4}&I_{3\times 3}\\ 0_{4\times 4}&0_{3\times 3}\end{bmatrix},G=0,W=\begin{bmatrix}I_{4\times 4}&0_{4\times 3}\\ 0_{3\times 4}&0.01I_{3\times 3}\end{bmatrix}$.
For the measurement model, the detections are converted into $\mathbf{z}=[u,v,s,r]^{\top}$, and we have
$H=\begin{bmatrix}I_{4\times 4}&0_{4\times 3}\end{bmatrix},\;V=\begin{bmatrix}I_{2\times 2}&0_{2\times 2}\\ 0_{2\times 2}&10I_{2\times 2}\end{bmatrix}$